OpenStudy (summersnow8):

What is the pH of each of the following solutions? a) 5.0 x 10^-4 M HCL b) 5.0 x 10^-4 M KOH c) 0.15 M HCL d) 1.5 x 10^-2 M NaOH @photon336

OpenStudy (whpalmer4):

First thing you need to know is the hydronium ion concentration of the solution, which you can find from the concentration of the acid or base in that solution. To do that, you need to know whether they are strong (fully dissociate in sufficient water) or weak. Do you know that for the three listed here?

OpenStudy (summersnow8):

HCl = strong acid KOH = strong base NaOH = strong base

OpenStudy (whpalmer4):

Okay, good. That makes things easier. Each molecule of those will dissociate and give us an equivalent concentration of either H+ for the acid or OH- for base. We plug that into the pH formula to get our pH

OpenStudy (summersnow8):

so when they completely dissociate their concentrations with be the same for [H] and [OH]

OpenStudy (whpalmer4):

so, if our strong acid is 0.1 M HCl, for example, our [H3O+] value will be 0.1, and our pH can be calculated by \[pH = -\log_{10} [H_3O^+] = -\log_{10} 0.1 = 1\]

OpenStudy (summersnow8):

ahhh okkay

OpenStudy (whpalmer4):

for the bases, we can use \[pOH = -\log_{10} [OH^-]\] and \[pH = 14 - pOH\]

OpenStudy (summersnow8):

so for a. \[pOH = -\log_{10} \left[ 5.0 \times 10^{-4} \right] \] \[pOH = 3.30\] pH = 14 - pOH pH = 14 - 3.30 pH = 10.7

OpenStudy (whpalmer4):

That's my understanding, yes.

OpenStudy (whpalmer4):

that's b), not a)...

OpenStudy (summersnow8):

a would be 3.30

OpenStudy (whpalmer4):

yes

OpenStudy (summersnow8):

c. would be \[pH = -\log_{10} \left[ .15 M \right] \] pH =0.82 ?

OpenStudy (summersnow8):

d. pOH=−log[1.5×10^-2] pOH=1.82 pH = 14 - pOH pH = 14 - 1.82 pH = 12.18

OpenStudy (whpalmer4):

log base 10 of a number is the power to which 10 must be raised to get it. \[10^{-1} = 0.1\]so \[\log_{10} 0.1 = -1\]right? That means our log should be somewhere in that ballpark, and after we take the negative of the value of the log, we should have a positive number for the pH, yes?

OpenStudy (whpalmer4):

(I'm talking about c) here)

OpenStudy (summersnow8):

how would i calculate that with my calculator? what would I plug in

OpenStudy (summersnow8):

10^.15

OpenStudy (whpalmer4):

sorry, I misread your answer, I read the = as a -, my bad!

OpenStudy (whpalmer4):

you'll be old someday, too :-)

OpenStudy (summersnow8):

so, so far I have solved a, b, and d, but c is wrong still?

OpenStudy (whpalmer4):

the answer for c is pH 0.82 as you said. I just misread it and thought you had given -0.82 I agree with all of your answers.

OpenStudy (summersnow8):

yay!